Roy decides to buy a personal residence and goes to the bank for a $150,000 loan. The bank tells him that he can borrow the funds at 4% if his father will guarantee the debt. Roy's father, Hal, owns a $150,000 CD currently yielding 3.5%. The Federal rate is 3%. Hal agrees to either of the following:

  • Roy borrows from the bank with Hal's guarantee to the bank.
  • Hal cashes in the CD (with no penalty) and lends Roy the funds at 2% interest.

Hal is in the 32% marginal tax bracket. Roy, whose only source of income is his salary, is in the 12% marginal tax bracket. The interest Roy pays on the mortgage will be deductible by him.

Considering only the tax consequences, answer the following.

a. The loan guarantee:

Hal's interest income from the CDs would be $ before taxes and $ after taxes.

Roy's interest expense from the bank loan would be $ before taxes and $ after taxes.

This arrangement would produce an overall (Negative or Positive) cash flow after taxes to the family of $.

b. The loan from Hal to Roy:
Hal's tax on the imputed interest income from the loan to Roy would be $.

Roy's tax benefit from the imputed interest expense from Hal's loan would be $.

This arrangement would produce an overall (Negative or Positive)  cash flow after taxes to the family of $.

c. Which option will maximize the family's after-tax wealth? (The Loan from Hal to Roy or The Loan Guarantee)
Answer #1

{ Note - Question in solve in paragraph with calculation for your better understanding , please fill bracket self }

a.] The Loan Guarantee arrangement

• For Hal's Interest Income from Current Deposits

($150000*3.5%)=$5250 Before Tax

($150000*3.5%)-($150000*3.5%*32%)-(5250*3%)=3728 After Taxes.

• For Roy's Interest Expenses

($150000*4%)= $6000 Before Tax After Taxes.

($150000*4%)-($150000*4%*12%)= 5280 after Taxes.

• Net Negative Outflow of Family after Taxes

5280- 3728= $1553

Note : Assuming the Federal Tax paid on Interest Income is available for deduction.

b.] Loan From Hal to Roy

•For Hal's Tax on Imputed Interest from Roy

($150000*2%)- ($150000*2%*32%)= 2040

•Loss of Interest from Current Deposits = $3728

•Outflow is $1688

•For Roy's Interest on Loan from Hal

•Interest Outflow = ($150000*2)- ($150000*2%*12)= $2640

•Family Outflow =$2640+$1688= $4328

Note: The interest income Hal would have earned from Current deposit is also considered as his deemed loss.

c] . First Case of Bank Guarantee arrangement is considered suitable as it would generate lower outflow for family..

I am so glad you asked for help when you needed it. ?